If price declines from $450 to $350 and, as a result, quantity demanded increases from 1200 to 1500, what is the price elasticity of demand?
1
Expert's answer
2014-11-17T12:40:25-0500
Price elasticity of demand (PED or Ed) is a measure used in economics to show the responsiveness, or elasticity, of the quantity demanded of a good or service to a change in its price. More precisely, it gives the percentage change in quantity demanded in response to a one percent change in price (ceteris paribus, i.e. holding constant all the other determinants of demand, such as income).
If price declines from $450 to $350 and, as a result, quantity demanded increases from 1200 to 1500, the price elasticity of demand is:
Ed = (450 + 350)/(1200 + 1500)*(1500 - 1200)/(350 - 450) = 800/2700*300/(-100) = -8/9, so the demand is inelastic (|Ed| < 1).